The solubility of potassium chloride at 20 degree celsius is 34.7g in 100g of water.The density of the solution is 1.3g/ml.What is the w/v percentage of potassium chloride in the solution

Grade:12

2 Answers

Divyansh Singhvi
10 years ago

(34.7g/100g)*1.3g/ml=0.45

Kaushik
10 years ago

let volume of the solution be v

density=(34.7+100)/v = 1.3

134.7=1.3v

v=134.7/1.3 ml

%w/v=mass of KCl in 100 ml of solution.

134.7/1.3 ml contains 34.7 g of KCl.

100 ml contains:(100 x 34.7 x 1.3)/134.7=33.49%